### v1

# Can Carbon-Aware Electric Load Shifting Reduce Emissions? An Equilibrium-Based Analysis

Abstract
An increasing number of electric loads, such as hydrogen producers or data centers, can be characterized as carbon-sensitive, meaning that they are willing to adapt the timing and/or location of their electricity usage in order to minimize carbon footprints. However, the emission reduction efforts of these carbon-sensitive loads rely on carbon intensity information such as average carbon emissions, and it is unclear whether load shifting based on these signals effectively reduces carbon emissions. To address this open question, we investigate the impact of carbon-sensitive consumers using equilibrium analysis. Specifically, we expand the commonly used equilibrium model for electricity market clearing to include carbon-sensitive consumers that adapt their consumption based on an average carbon intensity signal. This analysis represents an idealized situation for carbon-sensitive loads, where their carbon preferences are reflected directly in the market clearing, and contrasts with current practice where carbon intensity signals only become known to consumers aposteriori (i.e. after the market has already been cleared). We include both illustrative examples and larger numerical simulations, including benchmarking with other methods, to illuminate the contributions and limitations of carbon-sensitive loads in power system emission reductions.
Index Terms:
Carbon reduction, equilibrium analysis, carbon emission intensity signal design
I Introduction
The Paris Agreement recognizes emissions abatement as a critical strategy for mitigating global warming. Due to both the significant emissions from electricity generation and the potential for reducing these emissions by increasing renewable and low-carbon generation technologies, the electricity sector has been at the forefront of implementing decarbonization policies. However, the inherent variability of renewable generation, which is highly dependent on weather conditions (e.g., wind speed, solar radiation), necessitates aligning electricity demand with supply availability to maintain reliability. Demand-side flexibility is a crucial component of this alignment, i.e. consumers must be incentivized to adjust their consumption in response to renewable generation surpluses or deficits. At the same time, a growing number of carbon-sensitive loads, such as data centers and hydrogen production facilities, are emerging. These entities actively optimize the timing and location of their electricity usage to reduce carbon footprints, driven either by pledges to environmental responsibility or economic incentives such as production tax credits. However, many of these consumers are unwilling to give up control of their power consumption e.g. by participating in a market, and instead prefer to locally optimize their consumption based on externally provided information such as carbon intensity signals. In this paper, we try to explore the coordination of carbon-sensitive loads with the rest of the electricity market clearing through a novel equilibrium formulation for an electricity market clearing where we model the impact of average carbon emissions on consumers’ decision making.
Whereas calculating and potentially penalizing carbon emissions from generation is a relatively straightforward task, the question of how to allocate emissions from a group of generators across a group of consumers is less obvious and subject to ongoing debate [1]. A common approach is to define a carbon intensity signal (with the most commonly used and widely available signal being average carbon emissions [1]) which is computed based on the average rate of emissions per megawatt-hour (MWh) of electricity consumed [2, 3]. To differentiate the impacts of locations on the carbon reduction, some researchers have advocated for the use of locational marginal carbon emissions [4, 5, 6, 7, 8], which provides information about how a change in the electricity consumption at a given location would impact total emissions. Other, more recently proposed metrics include locational average carbon emissions [9], which is based on the concept of carbon flow [10, 11], and adjusted locational marginal emissions [8], which combines characteristics of locational and average emissions.
While carbon intensity signals are generally used for carbon accounting, e.g. as outlined in [1], carbon-sensitive loads also leverage these signals to decide where and/or when to consume electricity. In practice, such adjustment would happen in a sequential manner, where the electricity market is first cleared assuming a given load profile and carbon intensity signals are then calculated based on the market outcome. Subsequently, carbon-sensitive consumers utilize these carbon intensity signals to adapt their consumption across time and space, with the goal of achieving carbon reductions. The impact of these changes have been studied in the context of real-time market clearing [7, 12], where the updated load profile becomes the input to the next market clearing, and in the context of day-ahead market clearing [8], where the updated load profile is taken into account in the intra-day market. As observed in [7, 12, 8], this sequential approach leads to suboptimal outcomes because consumers act on carbon intensity signals arising from the previous market clearing, without insight into how their (and other consumers/generators) actions may change the generation dispatch and potentially impact the carbon intensity signals in the next market clearing.
Our recent work in [13] proposes a novel electricity market clearing model that incorporates the allocation of emissions from generations to loads and allows for consideration of consumer-side carbon costs. Continuing to further this line of research, in this paper, we propose a game-theoretic model, where all market participants jointly identify a solution to the market clearing by solving an equilibrium problem that includes average carbon signal and their impacts on electricity consumption. As we mentioned before, there are several different methods to define the carbon emission intensity of electricity consumption. We choose to use average carbon signals because this metric is widely available, both from system operators and companies [14, 15], and because it is commonly used for carbon accounting under the Greenhouse Gas Protocol [16, 1].
The major contributions of this paper are twofold. First, we propose a novel equilibrium model considering the average carbon signal, which assumes that consumers have “perfect” knowledge of their own carbon emission intensity. This formulation provides a new perspective to evaluate the effectiveness of different carbon signals and contributes to new carbon signal design. To the best of our knowledge, this is the first paper that investigates the impact of carbon-sensitive consumers on carbon reductions from a game theoretic perspective. Second, we demonstrate how carbon-sensitive consumers behave in the equilibrium model against the sequential model. Based on the real-world IEEE RTS-GMLC system, we then compare the equilibrium model with benchmark models with or without considering the impact of carbon emissions.
The remainder of the paper is organized as follows: Section II describes our proposed equilibrium model. In Section III, we show results and conclusions from our numerical study, while Section IV summarizes and concludes.
II Carbon-aware Equilibrium Electricity Market Clearing Model
In this section, we introduce the proposed equilibrium problem formulation. After that, we employ a simplified three-bus system to illustrate how loads adapting to carbon signals impact the generation dispatch and carbon emissions of the overall power system.
However, before diving into the details of our formulation, we first introduce some basic background on equilibrium models in the electric sector. Equilibrium models, or game-theoretic models, are commonly employed to investigate strategic behavior in deregulated electricity markets [17, 18]. In such models, there is typically three (in the electricity pool model, which ignores transmission constraints) or four (in the transmission-constrained model) types of participants, namely generators, consumers, transmission owners, and ISO. In the traditional carbon-agnostic electricity market, generators, consumers, and transmission owners aim at maximizing their profits given electricity prices as input parameters, and the ISO determines the market clearing price given power generation and electricity consumption as input parameters. All participants are, between them, playing a noncooperative game, and thus the optimal solution to this game is defined as a Nash equilibrium, corresponding to a situation where no participant can improve their outcomes by unilaterally changing their decisions. Under a price-taking assumption (no strategic bidding or market power) the Nash equilibrium problem is equivalent to a standard market clearing optimization problem, i.e., solutions from the equilibrium problem are the same as those from the optimization problem. The equilibrium can often be rewritten as a complementarity problem; more details, examples and formulations can be found in [19, 20, 21].
II-A Model Formulation
We next introduce our proposed equilibrium formulation with carbon-sensitive loads. As in the typical equilibrium model for electricity markets, we consider generators, consumers, transmission owners, and the ISO as participants. As is common, generators determine their optimal power generation based on electricity prices and the transmission owner maximizes their profit by buying power at a bus and selling it back at another one. Different from the standard model, consumers determine their electricity consumption considering both electricity prices and carbon emissions, and the ISO determines both the price and the average carbon signal. Below, we provide the parametric optimization problems solved by the generators, consumers, and transmission owners, respectively, and the complementarity conditions involving the ISO.
We consider an electric power network with the set of nodes, loads, transmission lines, and generators denoted $\mathcal{N}$, $\mathcal{D}$, $\mathcal{L}$ and $\mathcal{G}$, respectively. Let $\mathcal{G}_{i}\subset\mathcal{G}$ and $\mathcal{D}_{i}\subset\mathcal{D}$ be the subset of generators and loads connected to the node $i$, and $(i,j)\in\mathcal{L}$ denote the transmission line from node $i\in\mathcal{N}$ to node $j\in\mathcal{N}$. Let $g\in\mathcal{G}$ and $d\in\mathcal{D}$ denote the indexes for generators and consumers, respectively.
Generators: Similar to the conventional electricity market, each generator in our model aims to maximize profit under generation constraints. Specifically, given the electricity price $p_{i}$ as input, the generator solves the following optimization problem to determine the optimal generation output that maximizes profits,
$\max_{P_{G,g}}$ $(p_{i:g\in\mathcal{G}_{i}}-c_{G,g})\cdot P_{G,g}$ (1a)
$s.t.$ $P_{G,g}^{\min}\leq P_{G,g}\leq P_{G,g}^{\max},$ (1b)
where $P_{G,g}$ is the optimal generation output (decision variables), $c_{G,g}$ is the generation cost for generator $g$, $p_{i:g\in\mathcal{G}_{i}}$ represents that the electricity price at the node $i$ to which the generator $g$ is connected, and the constraint (1b) ensures that the chosen generation is feasible with $P_{G,g}^{\min}$ and $P_{G,g}^{\max}$ being generation limits.
Consumers: Each consumer in our model maximizes their benefits of consuming electricity, but different from a standard formulation, the consumers consider both price and carbon emissions in their decisions. This gives rise to a carbon-dependent consumer utility, which we define as $r_{D,d}-\lambda\cdot c_{D,d}$. Here, $r_{D,d}$ is the revenue for consumers derived from consuming electricity, $c_{D,d}$ denotes the carbon cost, given in units of [\$/tons $CO_{2}$] and $\lambda$ is the value of the average carbon signal. Specifically, given price $p_{i}$ and the average carbon signal $\lambda$ calculated by ISO in the later part, each consumer solves the following problem:
$\max_{P_{D,d}}$ $(r_{D,d}-p_{i:d\in\mathcal{D}_{i}}-\lambda\cdot c_{D,d})\cdot P_{D,d}$ (2a)
$s.t.$ $P_{D,d}^{\min}\leq P_{D,d}\leq P_{D,d}^{\max},$ (2b)
where $p_{i:d\in\mathcal{D}_{i}}$ represents that the electricity price at the node $i$ to which the consumer $d$ is connected. Each consumer incorporates the carbon term $\lambda\cdot c_{D,d}$ into their utility evaluation. In practice, the cost parameter $c_{D,d}$ may be directly tied to concrete costs, such as carbon emission penalties from carbon taxes or cap-and-trade schemes, or could be an internally set “carbon cost”, reflecting how much revenue the consumer is willing to forgo to avoid carbon emissions. Note that besides using carbon-dependent utilities to quantity the impacts of carbon emissions on consumers’ decision-making, there are also other ways to quantify this impact, e.g., setting carbon emission limits [22, 23]. However, it can be very challenging for consumers to define what the values of the carbon emission cap should be.
Transmission Owners: Transmission owners act as spatial arbitragers and aim to maximize the power transmission profit given electricity prices $p_{i}$. Specifically, the transmission owner maximizes their profit by buying power at a bus and selling it back at another one (arbitrage through electricity price difference between two bus nodes), which is characterized by the following mathematical optimization problem:
$\max_{\theta_{i}}$ $\sum_{i}p_{i}\cdot\left(\sum_{j,(i,j)\in\mathcal{L}}\beta_{ij}(\theta_{j}-\theta_{i})\right)$ (3a)
$s.t.$ $-F_{ij}^{\rm{lim}}\leq\beta_{ij}(\theta_{i}-\theta_{j})\leq F_{ij}^{\rm{lim}},\quad\forall(i,j)\in\mathcal{L},$ (3b)
$\theta_{ref}=0,$ (3c)
where $\beta_{ij}\in\mathbb{R}$ denoting the susceptance value of the transmission line $(i,j)$ from node $i$ to node $j$. Constraints (3b) are the transmission line limits, where $F_{ij}^{\rm{lim}}$ represents the transmission capacity (which we assume is the same in both directions). Constraint (3c) sets the voltage angle at the reference node to zero.
ISO: The ISO clears the market by determining the electricity trading price at each node and provides the average carbon signal. Specifically, they set each nodal electricity price $p_{i}$ such that the supply-demand power balance
$\sum_{d\in\mathcal{D}_{i}}P_{D,d}+\sum_{j:(i,j)\in\mathcal{L}}\beta_{ij}(\theta_{i}-\theta_{j})=\sum_{g\in\mathcal{G}_{i}}P_{G,g}$ (4)
holds. The average carbon signal $\lambda$ is defined as the total carbon emission divided by total demand, which can be written as
$\lambda\sum_{d\in\mathcal{D}}P_{D,d}=\sum_{g\in\mathcal{G}}e_{G,g}\cdot P_{G,g},$ (5)
where $e_{G}=\{e_{G,g}:\forall g\in\mathcal{G}\}$ is the carbon emission intensity vector for all generators (for generators using renewable sources, the carbon emission intensity values are zero).
Note that (4) and (5) are written here as equality constraints but form a square complementarity system when paired with $p$ and $\lambda$, respectively. The number of equations in (4) matches the size of the variable $p$; similarly with (5) and $\lambda$. Both $p$ and $\lambda$ are parameters in the optimization problems (1)–(3).
The overall Nash equilibrium problem consists of the collection of parametric optimization problems (1), (2), (3) and the parametric complementarity conditions (4) and (5).
Remark 2: When all $c_{D}=0$, our equilibrium problem reduces to the standard social welfare maximization problem. The equilibrium framework seeks to coordinate all participants toward a consistent set of values for $P_{G}$, $P_{D}$, $p$, and $\lambda$ that simultaneously ensures power balance and aligns with the optimal solutions of each participant’s individual optimization problem. Unlike the sequential method, where inaccuracies may arise from relying on lagged market-clearing signals (see [8, 7, 12]), the equilibrium method uses synchronized decision-making based on perfect real-time carbon signals, which we believe is better for analyzing the effectiveness of different carbon signals. We note that, similar to the standard market-clearing optimization problem, the equilibrium problem may have multiple solutions (e.g., when multiple generators at a node share identical costs).
II-B Three-bus Example Illustration
In order to illustrate the outcomes of using an equilibrium model with carbon-dependent utility for consumers, we employ a simple three-bus system adapted from Example 6.2.2 in [20], as seen in Fig. 2. The network includes three generators and three consumers, with one generator and consumer located at each bus. Since the carbon intensity of the generators is not included in the original three-bus system data, we define two carbon emission intensity cases to see different results of load shifting among consumers in Table I. In Case I, we assume that higher cost generators have lower carbon emissions. Specifically, the most expensive generator, located at Bus 2, has the lowest carbon intensity, while the cheapest generator, located at Bus 3, has the highest carbon intensity. The medium cost generator at Bus 1 has an intermediate carbon intensity. In Case II, we assume that higher cost generators have higher carbon emissions. We assume that all consumers have a carbon cost of $c_{D,d}=\$20/$tons.
TABLE I: Parameters of Generators and Consumers.
Bus (#) 1 2 3
Consumers $P_{D}^{\rm{min}}$(MW) 4 16 12
$P_{D}^{\rm{max}}$(MW) 6 24 18
$r_{D}$(\$/MWh) 18 20 21
Generators $P_{g}^{\rm{min}}$(MW) 0 0 0
$P_{G}^{\rm{max}}$(MW) 20 10 25
$c_{G}$(\$/MWh) 8 10 6
Case I $e_{G}$(tons/MWh) 0.6 0.2 1
Case II 0.6 1 0.2
(a) Case I. Higher cost generators have lower carbon intensity.
(b) Case II. Higher cost generators have higher carbon intensity.
Figure 1: Results for different generation cost-carbon intensity pairs.
(a) Case I.
(b) Case II.
Figure 2: Generation and consumption for Case I and Case II in Table I.
II-B1 Electricity Pool Case
We first consider the electricity pool case, which omits consideration of transmission constraints. The results for this case are shown in Fig. 1 (a) and (b) and detailed generation dispatch (purple numbers) and demand values (green numbers) are shown in Fig. 2. Fig. 1 (a) shows results for Case I, with the black line showing the generation cost curve (with generators ordered from the cheapest to the most expensive) and the purple line showing the total carbon emission curve for Case I. The total carbon emissions $E_{tot}$ is calculated by multiplying the amount of generation with the corresponding carbon intensities, i.e., $E_{tot}=\sum_{g\in\mathcal{G}}e_{G,g}\cdot P_{G,g}$ based on the generation dispatch order. The red and purple points on the black generation cost curve represent the dispatch results for two cases, respectively. The purple shadowed areas highlight the effects of our proposed model compared with the case without considering carbon emissions. Similarly, Fig. 1 (b) shows results for Case II, with the black line showing the generation cost curve (which is the same as for Case I) and the green line showing the total emission curve for Case II.
Our first observation from Fig. 1, is that the only impact of considering the carbon-dependent utility is that the consumers are less willing to pay for electricity with a non-zero average carbon footprint, and hence reduce their electricity demand. The generation merit order (represented by the generator cost curve) creates an ordering of the generators. The demand reduction sequentially impacts the generators, starting with the marginal one which reduces production, and if needed, continues in descending order. If the affected generators have non-zero emissions, reductions in demand have the intended effect of decreasing emissions, as witnessed in both Case I and Case II. However, the fact that carbon-aware loads can only sequentially impact the generators according to the aforementioned ordering, which is not related to the ranking of generators by their emissions, is a fundamental limit of carbon-aware load shifting in our equilibrium model.
Our second observation from Fig. 1 is that the demand reduction and associated reduction in emissions is very different between Case I and II. In Case I, the generation dispatch includes low-cost, carbon-intensive generation which leads to a high average carbon value $\lambda=0.913$. Since $\lambda$ is high, the carbon-dependent consumer utility is much lower than the normal consumer utility, leading to a large reduction in demand from 48 to 32 MW, or 33%, when considering carbon. However, since the marginal generator has low emissions, this reduction in demand only decreases emissions from 37 to 29.2 tons, or 21%. This suggests that each unit of demand reduction achieves a lower-than-average impact on emissions in Case I. Conversely, in Case II, the generation dispatch includes low-cost, low-carbon generation which leads to a low average carbon value $\lambda=0.4$. With a lower $\lambda$, the carbon-dependent consumer utility is closer to the normal consumer utility, and the demand is only reduced to 46.7 MW, or less than 3%, when considering carbon. This reduction in demand reduces emissions from 20 to 18.7 tons, or 6.5%. Thus, in Case II, each unit of demand reduction achieves a higher-than-average impact on emissions, but the overall demand and carbon reductions are small.
This example highlights why the average carbon signal may not be the best choice, as it does not provide any carbon intensity information about the marginal generator, which is the only generator impacted by the load shift, and thus does not accurately predict how a change in demand may increase or decrease emissions. Reducing demand in a situation with low average carbon emissions, such as Case II, can have a greater impact on emissions than reducing demand in situations with high average emissions, such as Case I.
(a) Base case ($c_{D,d}=0$ for $\forall d\in\mathcal{D}$).
(b) Case I. Higher cost-lower carbon generators.
(c) Case II. Higher cost-higher carbon generators.
Figure 3: Results comparison for transmission-constrained cases.
II-B2 Transmission-constrained Case
To explain the impacts of transmission line constraints in our equilibrium model, we reduce the transmission capacity on the transmission line from Bus $2$ to Bus $3$ so it becomes congested. This congestion increases power generation at Bus 2 and decreases power generation at Bus 1. Fig. 3 shows the dispatch for the base case ($c_{D}=0$) and Case I and Case II ($c_{D}=20$), all with transmission line limits. For Case I, the shift of generation from generator 1 to 2 increases the average carbon emissions from $\lambda=0.608$ (for electricity pool case) to $\lambda=0.767$ in the base transmission-constrained case (Case I in Fig. 3a). Introducing carbon-sensitive loads reduces the load from 48 to 32 MW, as in the case without transmission constraints. However, due to congestion, both generator 1 and 2 are now marginal and are impacted by the demand reduction leading to a slightly higher reduction in emissions of 22.5%. Nevertheless, the average emissions increase to $\lambda=0.89$ after the load reduction. For Case II, shifting generation from generator 1 to 2 reduces the average carbon emissions to $\lambda=0.433$. This $\lambda$ is low enough to allow consumers to continue to consume their maximum load, and as a result, the generation dispatch does not change after we introduce carbon-sensitive loads. This is despite the fact that the two marginal generators in this case have higher emissions than in Case I.
This example highlights that transmission constraints impact the marginal generators serving a carbon-sensitive loads, and thus impacts the effectiveness of demand reductions from carbon-sensitive loads.
III Numerical Studies
This section conducts extensive numerical studies to demonstrate how equilibrium method could impact electricity market clearing. All the tests are performed on a laptop computer with an Apple M3 Pro CPU and 36GB RAM. The equilibrium problem is solved using EMP [24]. We adopt the IEEE RTS-GMLC system [25] for numerical analysis. This system has $73$ buses, $158$ generators, and $120$ lines. For generator fuel types, this system covers natural gas, oil, coal, and renewable sources (solar, wind, and hydro). The emission intensities of different generators are assigned according to the US Department of Energy [26] and are listed in Table II. For simplicity, we consider each non-zero load in the system as a consumer, leading to a total of $51$ consumers located at different buses. According to [27], carbon prices in the world-wide emission trading markets range from $10$ to $100$\$/tons. Based on these statistics, we assume that consumers will have carbon prices $c_{D}$ in the range between $[10,100]$ in the following experiments. For consumers’ demand flexibility, we use the values in the IEEE RTS-GMLC system as the maximum values and 80% of these values as the minimum values.
TABLE II: Generator carbon emission intensities.
Type Carbon Emission Intensity(tons/MWh)
Natural Gas 0.6042
Oil 0.7434
Coal 0.9606
Wind, Solar, Hydro 0
III-A Evaluation Metrics and Benchmark Formulations
In our case study, we consider the following quantities when evaluating market clearing solutions:
  • Total generation: $\sum_{g\in\mathcal{G}}P_{G,g}$;
  • Total generation cost: $\sum_{g\in\mathcal{G}}c_{G,g}P_{G,g}$;
  • Total carbon emissions: $\sum_{g\in\mathcal{G}}e_{G,g}P_{G,g}$;
  • Average carbon emissions: $\lambda$.
We consider three benchmark models for our analysis:
  • Market Clearing with Fixed Demands (MCFD): The traditional (carbon-agnostic) electricity market solves the market clearing problem assuming a fixed demand. For simplicity, we assume that the demand of each consumer is fixed at their maximum values $P_{D,d}^{\max}$.
  • Market Clearing with Demand Flexibility (MCDF): In this case, consumers have flexibility in their electricity consumption, but do not consider carbon costs. Note that this model is equivalent to a version of the proposed equilibrium model where all consumers set carbon costs $c_{D}=0$.
  • Sequential Method: We consider the sequential method first proposed for to represent geographical load shifting in [28] and extended to consider multiple timesteps in [8]. This method consists of three steps:
    Step $1$: We clear the day-ahead electricity market, without considering consumer carbon costs. This step is equivalent to solving the MCFD model.
    Step $2$: We calculate the average carbon emissions from the market clearing results, namely $\lambda_{b}$. We then solve (2) with the calculated average carbon emission to represent how carbon-sensitive loads would adapt their load.
    Step $3$: Finally, we solve another market clearing similar to the MCDF, representing the intra-day electricity market, with updated load profiles. We denote the average carbon emission based on these market clearing results as $\lambda_{a}$.
III-B Analysis for Different Carbon Costs
We first analyze the impact of different carbon costs on the equilibrium results, assuming that all consumers are willing to submit non-zero carbon costs to help reduce carbon emissions. We compare cases where the carbon costs of consumers are randomly drawn from a uniform distribution with ranges $[10,40]$, $[30,60]$, $[50,80]$, and $[80,100]$, respectively. Table III shows the results. The top rows represent the two carbon-agnostic formulations (MCFD and MCDF), and the average carbon cost increases as we go further down in the rows.
TABLE III: The impact of different carbon costs on generation dispatch, system emissions, and other components in equilibrium models.
Cases
Total Generation
 [MWh]
Total Generation
Cost [\$]
Total Carbon
 [tons]
Average Carbon
 [tons/MWh]
$u_{d}^{\intercal}P_{D}$ [\$] $\lambda\sum\limits_{d\in\mathcal{D}}c_{D,d}\cdot P_{D,d}$ [\$]
MCFD 8550 63748 3001.8 0.351 457039 0
MCDF 8550 63748 3001.8 0.351 457039 0
$[10,40]$ 8550 63748 3001.8 0.351 457039 80271.3
$[30,60]$ 8515.8 63358 2969 0.349 456301.4 138965.4
$[50,80]$ 8368.4 61677.3 2827.4 0.338 452348 189575.5
$[80,100]$ 8159.2 59986.7 2680.8 0.329 445700.2 244915.1
We first observe that the two carbon-agnostic market clearing models (MCFD and MCDF) lead to the same solutions. This indicates that the consumer utility values $r_{D}$ are high enough to cause loads to consume at their highest level in the formulation with load flexibility, but no carbon costs. Next, we consider the impact of low carbon costs in the range $[10,40]$ (the third row of Table III). For these carbon costs, we observe a negligible impact on total power generation, indicating that the loads are still consuming the same amount of electricity despite occurring a carbon cost. We conclude that when carbon costs are low, they do not impact consumers’ electricity consumption. Further, we consider the case with higher carbon costs in the ranges $[30,60]$, $[50,80]$, or $[80,100]$ (lower three rows of Table III). As the carbon costs increase to these levels, the total demand decreases, and the total generation decreases accordingly. This happens because some loads choose to consume less, as the carbon-dependent utility is too low (because the carbon cost becomes too high). As the total generation decreases, the total generation cost, total carbon emissions, and average carbon emissions also decrease.
III-C Comparing against the Sequential Method
We next seek to compare our results to those achieved if we use the sequential method. We run the sequential method with the same input data as we provided to the equilibrium model in Section III-B. The results for both methods under different carbon cost ranges are shown in Fig. 4.
(a) Total Generation (MWh).
(b) Total Generation Cost (\$).
(c) Total Carbon (tons).
(d) Average Carbon (tons/MWh).
(e) Total Carbon Cost (\$).
Figure 4: Comparison of equilibrium model with sequential load shifting model.
We first focus on the case with carbon costs in $[10,40]$. Due to carbon costs being not high enough, both the equilibrium method and the sequential method share the same optimal solutions compared to carbon-agnostic results. With higher carbon costs in the ranges at or above $[30,60]$, the total generation is reduced due to demand reductions. We observe that the results from the equilibrium method and the sequential method are the same for carbon costs of $[30,60]$ and $[80,100]$, but different for the case with $[50,80]$. Specifically, the sequential method has lower total generation (indicating larger demand reduction) and lower average carbon emissions as compared with the equilibrium method. We next explain these differences by comparing the average carbon emissions observed by the sequential method when performing load shifting in step $2$ (which are based on the initial market clearing) with the average carbon emissions after the market has cleared again in step $3$, as well as the average carbon emissions from the equilibrium method.
For the [50,80] case, consumers in the sequential model observe an average carbon emission of $\lambda_{b}=0.351$ after the initial market clearing (Step $1$). This $\lambda_{b}$ is high enough for some loads to reduce their consumption from $P_{D,d}^{\max}$ to $P_{D,d}^{\min}$. This leads to a reduction in load which lowers the average carbon emission to $\lambda_{a}=0.336$ after the market clears again (Step $3$). This $\lambda_{a}$ is low enough for some loads to want to increase their consumption again; however, they do not know this at the time of shifting and are unable to act after the market has cleared the second time (i.e., consumers conduct load shifting in Step 2 only using last clearing results but without knowing the impacts of their behaviors on the next clearing results). With the equilibrium model, the loads can “see” the impact of their demand reductions when the market is cleared. In this case, the market participants “agree” to a dispatch where the average carbon emissions $\lambda=0.338$. With this average carbon emissions, more loads choose to consume $P_{D,d}^{\max}$ as compared with the sequential load shifting case. This leads to a smaller load reduction, which corresponds to higher total emissions. However, the solution to the equilibrium model is more aligned with the objective of maximizing the (carbon-dependent) utility of the loads.
Even in the cases where we get similar solutions for the sequential and equilibrium cases, i.e. with carbon cost ranges of $[30,60]$ and $[80,100]$, there are differences between the average carbon emissions $\lambda_{b}$ seen by the loads when doing the shifting (in step 2) and the resulting $\lambda_{a}$ after the market has cleared again, even though the latter is exactly the same as the ones computed by the equilibrium model. For example, with carbon costs $[30,60]$, these two $\lambda$ values are $\lambda_{b}=0.351$ while performing the shifting and $\lambda_{a}=0.349$ after the subsequent market clearing. However, this change in $\lambda$ is too small to have a direct impact on the optimal amount of demand reduction. This is because our model exhibits a bang-bang behavior, where loads switch from $P_{D,d}^{\max}$ to $P_{D,d}^{\min}$ once the term $r_{D,d}-p_{i:d\in\mathcal{D}_{i}}-\lambda\cdot c_{D,d}$ goes from positive to negative. If the change in average carbon emissions is small, it might be that no load experiences this switch. In the case with $[30,60]$, the equilibrium model arrives at an optimal average carbon emission $\lambda=0.349$, and the demand reduction of the equilibrium model matches the demand reduction of the sequential model. This indicates that the change in average carbon emissions from $\lambda_{b}=0.351$ to $\lambda_{a}=0.349$ did not change the optimal consumption level of any load.
IV Conclusions
This paper proposes a novel equilibrium model to analyze the impact of carbon-sensitive loads on carbon emissions of power systems based on the average carbon signal. The equilibrium method provides a new perspective on how carbon-sensitive loads utilize their demand flexibility to consume cleaner electricity. Our case study demonstrates that the equilibrium method is equivalent to the sequential method when carbon costs of consumers are in high or low ranges. The equilibrium method provides a better carbon signal for consumers to make carbon-aware decision-making than the sequential method, where consumers may overreact or insufficiently react to the carbon signal computed from the last market clearing without knowing the impacts of their behaviors on the next market clearing results.
We adopt the average carbon signal to guide consumers’ behaviors for carbon reductions. Although it is easy to compute, it neglects the impacts of different locations, which has been proven important for carbon reductions in [29]. In future work, we plan to expand our analysis to consider locational carbon emission signals [9, 30, 11, 13].